OpenStudy (anonymous):

Point E is located at (–2, 2) and point F is located at (4, –6). What is the distance between points E and F? is it 10

OpenStudy (carniel):

(6, -8)

OpenStudy (carniel):

Just count the spaces from where -2,2 is all the way to 4,-6

OpenStudy (anonymous):

A.square root 52 B.square root 28 C.10 D.square root 20

OpenStudy (carniel):

o_O?

OpenStudy (anonymous):

I say10

OpenStudy (carniel):

Give me a sec

OpenStudy (anonymous):

yay i got it right thank you

OpenStudy (waynex):

It is 10. Subtract an x from one point from the other. Do the same with the y's. Square each one, and then add and take the square root. \[\sqrt{(-2-4)^2+(2-6)^2}\] Sorry, I adjusted the equation.

OpenStudy (waynex):

You're welcome

OpenStudy (anonymous):

Find the length of the hypotenuse of a right triangle with legs of 9 cm and 12 cm. is the answer 15cm

OpenStudy (carniel):

9 squared + 12 squared = 144+81= 225 and 15 squared = 225 so yes